OpenStudy (anonymous):

please help medals What word describes the slope of the line passing through points (5, 12) and (5, -13)

OpenStudy (anonymous):

@YanaSidlinskiy @Dark_Riddles @DangerousJesse

OpenStudy (anonymous):

A. Positive B. Zero C. Negative D. Undefined

OpenStudy (shinalcantara):

\[m = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] \[m = \frac{ -13 - 12 }{ 5-5 }\] \[m = \frac{ -25 }{ 0 }\] what do you call number with zero as the denominator?

OpenStudy (anonymous):

idk im not very smart

OpenStudy (shinalcantara):

when you solve it in calculator, you'll get math error which only means it is undefined

OpenStudy (anonymous):

thank you